Find the ratio of the optical power (force) of the lenses made from glass and

diamond with the same radius of curvature. The index of refraction of glass is ng = 1.5,

the index of refraction of diamond is nd = 2.42 and the index of refraction of air is n = 1.


1
Expert's answer
2022-01-10T13:57:19-0500

"d=\frac{\sqrt{n_d^2-n^2}}{\sqrt{n_g^2-n^2}}=\sqrt{\frac{2.42^2-1}{1.5^2-1}}=1.97."